# Consensus Complementarity Control for Multi-Contact MPC

### v2

@@ I Introduction
    We introduce a novel convex projection that can improve the solve times for the C3 algorithm by more than 3 times as shown in Section VII-B. We also include theoretical analysis of this approximate projection, connecting its limiting behavior to that of an LCP Projection.
@@ III-A Linear Complementarity Problem
[…]
@@
Here, the vector $\lambda$ typically represents the contact forces and slack variables ([39]), $y$ represents the gap function and orthogonality constraint embeds the hybrid structure.
@@ IV-B Anitescu Formulation
$v_{k+1}=v_{k}+\Delta tJ_{f}\begin{bmatrix}q_{k}\\ v_{k}\\ u_{k}\end{bmatrix}+D\lambda_{k}+d_{v}.$
@@ V-B1 Consensus Formulation
$\min_{z}\quad$ $c(z)+\mathcal{I}_{\mathcal{D}}(z)+\mathcal{I}_{\mathcal{C}}(z)+\sum_{k=0}^{N-1}\mathcal{I}_{\mathcal{H}_{k}}(\delta_{k})$ (16)
@@ V-B2 ADMM Steps for the Consensus Formulation
$\min_{\delta_{k}}\quad$ $(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))^{T}\rho G_{k}(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))$ (22)
@@ V-B3 Projection Operation
$\min_{\delta_{k},s_{k}}\quad$ $(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))^{T}U(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))$ (23)
@@
$\min_{\delta_{k}}\quad$ $\frac{1}{2}||\delta_{k}^{x}-x_{d}||_{Q_{x}}+\frac{1}{2}||\delta_{k}^{u}-u_{d}||_{Q_{u}}$ (24)
@@ VI Illustrative Examples
For these results, OSQP [52] is used to solve quadratic programs and Gurobi [53] is used for mixed integer programs. PATH [54] and Lemke’s algorihm have been used to solve LCP’s. SI units (meter, kilogram, second) are used. The experiments are done on a desktop computer with the processor Intel i7-11800H and 16GB RAM. Reported run-times include all steps in the algorithm. The code for all examples is available [^2] https://github.com/AlpAydinoglu/coptimal, and we optimized the code implementation leading to algorithm speeding up especially for the finger gaiting and pivoting examples compared to the preliminary conference version. We have added the specific systems matrices for all the examples333 https://github.com/AlpAydinoglu/coptimal/blob/main/system_matrices.txt . Experiments are also shown in the supplementary video444 https://youtu.be/XBzlyNhKl8w .
@@ VI-C Cart-pole with Soft Walls
We design a controller where $s=10$, $\rho=0.1$, $G_{k}=I$, $\rho_{s}=2$, $N=10$, and $\Delta t=0.01$ for $\mathcal{L}_{\Delta t}$. There is a clear trade-off between solve time and planning horizon [56]. We test all three projection methods described in Section V and report run-times (single solve of (23)) on Table III averaged for $800$ solves. We also report the average of cost-to-go value assuming all of the methods can run at $100$ Hz. Even though the MIQP projection is usually better at exploring new contacts, it does not always result in a better cost. We note that the controller can run slightly faster than $240$ Hz if the LCP-based projection is used. For this example, the QP in (21) has no inequality constraints and can therefore the KKT conditions can be directly solved.
@@ VII-A Simulation example: Trajectory tracking with a single ball
One might easily warm-start C3 with an initial trajectory, and corresponding nominal hybrid plan. Here, to demonstrate that the algorithm can reliably discover trajectories in real-time from scratch, we explicitly do not provide such a warm-start. Instead, we add a time-based heuristic as shown in Figure 10. In Phase I ($1$ s long), we pick the end-effector bias slightly above the ball’s current state (shown in green). In Phase II ($0.5$ s long), we pick a set of end-effector biases so that it ends up slightly behind the ball with respect to the desired ball location (shown in black) and also increase the cost on end-effector deviation. It is important to note that we do not specify when and how the contact interactions should happen. All the end-effector biases are above the ball, and algorithm decides that the end-effector should interact with the ball on its own (as there is also penalty on ball location error). Moreover, C3 can decide that an interaction should happen in either Phase I and Phase II, but we softly encourage interactions to happen in Phase I and the end-effector to reposition in Phase II.
@@ VII-B Simulation example: Trajectory tracking with multiple balls
where $x_{c}$ is a constant value and $d_{y}$ is the desired translation on y-direction. We also follow the same heuristic as the previous section and set the end-effector bias with respect to the ball that has made the least progress in the $y$ direction.
@@ VIII-A ADMM
$\min_{z}\quad$ $c(z)+\mathcal{I}_{\mathcal{D}}(z)+\mathcal{I}_{\mathcal{C}}(z)+\sum_{k=0}^{N-1}\mathcal{I}_{\mathcal{H}_{k}}(\delta_{k})$ (26)
@@ VIII-B Proof of Lemma 3
$\operatorname{error_{comp}}$ $=\delta_{k}^{\lambda}\varTheta=\left(\delta_{k}^{\lambda}-\mu+\mu\right)^{T}\varTheta=\left(\delta_{k}^{\lambda}-\mu\right)^{T}\varTheta$ (30)

### v3

@@ I Introduction
    We introduce a novel convex projection that can improve the solve times for the C3 algorithm by more than 3 times as shown in Section VII-B. We also include theoretical analysis of this approximate projection, connecting its limiting behavior to linear complementarity problems.
@@ III-A Linear Complementarity Problem
Given a vector $q\in\mathbb{R}^{m}$, and a matrix ${F\in\mathbb{R}^{m\times m}}$, the $\text{LCP}(q,F)$ describes the following mathematical program:
$\underset{}{\text{find}}$ $\lambda\in\mathbb{R}^{m}$
@@
Here, the vector $\lambda$ typically represents the contact forces and slack variables [39], $y$ represents the gap function and orthogonality constraint embeds the hybrid structure.
@@ IV-B Anitescu Formulation
$v_{k+1}=v_{k}+\Delta tJ_{f}\begin{bmatrix}q_{k}\\ v_{k}\\ u_{k}\end{bmatrix}+D\lambda_{k}+\Delta td_{v}.$
@@ V-B1 Consensus Formulation
$\min_{z}$ $c(z)+\mathcal{I}_{\mathcal{D}}(z)+\mathcal{I}_{\mathcal{C}}(z)+\sum_{k=0}^{N-1}\mathcal{I}_{\mathcal{H}_{k}}(\delta_{k})$ (16)
@@ V-B2 ADMM Steps for the Consensus Formulation
$\min_{\delta_{k}}$ $(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))^{T}\rho G_{k}(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))$ (22)
@@ V-B3 Projection Operation
$\min_{\delta_{k},s_{k}}$ $(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))^{T}U(\delta_{k}-(z_{k}^{i+1}+w^{i}_{k}))$ (23)
@@
$\min_{\delta_{k}}$ $\frac{1}{2}||\delta_{k}^{x}-x_{d}||_{Q_{x}}+\frac{1}{2}||\delta_{k}^{u}-u_{d}||_{Q_{u}}$ (24)
@@ VI Illustrative Examples
For these results, OSQP [52] is used to solve quadratic programs and Gurobi [53] is used for mixed integer programs. PATH [54] and Lemke’s algorihm have been used to solve LCPs. SI units (meter, kilogram, second) are used. The experiments are done on a desktop computer with the processor Intel i7-11800H and 16GB RAM. Reported run-times include all steps in the algorithm. The code for all examples is available [^2] https://github.com/AlpAydinoglu/coptimal, and we optimized the code implementation leading to algorithm speeding up especially for the finger gaiting and pivoting examples compared to the preliminary conference version. We have added the specific systems matrices for all the examples333 https://github.com/AlpAydinoglu/coptimal/blob/main/system_matrices.txt . Experiments are also shown in the supplementary video444 https://youtu.be/L57Jz3dPwO8 .
@@ VI-C Cart-pole with Soft Walls
We design a controller where $s=10$, $\rho=0.1$, $G_{k}=I$, $\rho_{s}=2$, $N=10$, and $\Delta t=0.01$ for $\mathcal{L}_{\Delta t}$. There is a clear trade-off between solve time and planning horizon [56]. We test three projection methods described in Section V and report run-times (single solve of (23)) on Table III averaged for $800$ solves. We also report the average of cost-to-go value assuming all of the methods can run at $100$ Hz. Even though the MIQP projection is usually better at exploring new contacts, it does not always result in a better cost. We note that the controller can run slightly faster than $240$ Hz if the LCP-based projection is used. For this example, the QP in (21) has no inequality constraints and can therefore the KKT conditions can be directly solved.
@@ VII-A Simulation example: Trajectory tracking with a single ball
One might easily warm-start C3 with an initial trajectory, and corresponding nominal hybrid plan. Here, to demonstrate that the algorithm can reliably discover trajectories in real-time from scratch, we explicitly do not provide such a warm-start. Instead, we add a time-based heuristic as shown in Figure 10. In Phase I ($1$ s long), we pick the end-effector bias slightly above the ball’s current state (shown in blue). In Phase II ($0.5$ s long), we pick a set of end-effector biases so that it ends up slightly behind the ball with respect to the desired ball location (shown in red) and also increase the cost on end-effector deviation. It is important to note that we do not specify when and how the contact interactions should happen. All the end-effector biases are above the ball, and algorithm decides that the end-effector should interact with the ball on its own (as there is also penalty on ball location error). Moreover, C3 can decide that an interaction should happen in either Phase I and Phase II, but we softly encourage interactions to happen in Phase I and the end-effector to reposition in Phase II.
@@ VII-B Simulation example: Trajectory tracking with multiple balls
where $x_{c}$ is a constant value and $d_{y}$ is the desired translation on $y$ direction. We also follow the same heuristic as the previous section and set the end-effector bias with respect to the ball that has made the least progress in the $y$ direction.
@@ VIII-A ADMM
$\min_{z}$ $c(z)+\mathcal{I}_{\mathcal{D}}(z)+\mathcal{I}_{\mathcal{C}}(z)+\sum_{k=0}^{N-1}\mathcal{I}_{\mathcal{H}_{k}}(\delta_{k})$ (26)
@@ VIII-B Proof of Lemma 3
$\operatorname{error_{comp}}$ $=(\delta_{k}^{\lambda})^{T}\varTheta=\left(\delta_{k}^{\lambda}-\mu+\mu\right)^{T}\varTheta=\left(\delta_{k}^{\lambda}-\mu\right)^{T}\varTheta$ (30)